OpenStudy (anonymous):

FIRST PERSON TO ANSWER THIS GETS A MEDAL AND A CHOCOLATE CHIP RASPBERRY COOKIE!!!! Complete the equation that is related to the given equation. x – (–6.9) = 42.8 x = 42.8 + OPTIONAL: LEAVE MAILING ADDRESS FOR COOKIE.

OpenStudy (anonymous):

x= 42.8-(-6.9)

geerky42 (geerky42):

:D

OpenStudy (anonymous):

49.7?

geerky42 (geerky42):

Yep

OpenStudy (anonymous):

Cookie shipping...

geerky42 (geerky42):

Yay can't wait to eat virtual cookie!

OpenStudy (anonymous):

Click link in email to eat.

jimthompson5910 (jim_thompson5910):

x – (–6.9) = 42.8 turns into x + 6.9 = 42.8 because subtracting a negative is the same as addition (think of the two negatives cancelling to get a positive)

OpenStudy (anonymous):

good ppl in this chat

jimthompson5910 (jim_thompson5910):

To solve x + 6.9 = 42.8, we subtract 6.9 from both sides put another way, we add -6.9 to both sides x + 6.9 = 42.8 x + 6.9 + (-6.9) = 42.8 + (-6.9) x+0 = 42.8 + (-6.9) x = 42.8 + (-6.9)

geerky42 (geerky42):

Thanks

OpenStudy (anonymous):

i yall were trees yall would be good trees

OpenStudy (anonymous):

if*

OpenStudy (anonymous):

Everybody gets one :)